OpenStudy (anonymous):

(5,3); m=3/5 using the point slope form y-y1=m(x-x1)

OpenStudy (shamil98):

Input your values then.. (5,3) --> (x,y) m = 3/5

OpenStudy (shamil98):

\[\Huge y - 3 = \frac{ 3(x-5) }{ 5 }\]

OpenStudy (anonymous):

are we still solving ?

OpenStudy (shamil98):

What does your problem ask you to do?

OpenStudy (anonymous):

to find the equation of the line

OpenStudy (shamil98):

Okay, so in slope-intercept form then?

OpenStudy (shamil98):

Or just point-slope form? because the equation i wrote is just in point-slope form.

OpenStudy (anonymous):

yeah and to solve the point slope form of the equation

OpenStudy (anonymous):

use the formula: y=mx+b y=3, x=5, m=3/5 3=3/5(5)+b 3=3+b 0=b =>y=3/5x